i need to solve this im stuck
use ln a=8 and ln b=3, to find each of the following
ln(ab)
ln(b/a)
ln(b^4)
ln √(a)

Answer by lwsshak3(11628)

use ln a=8 and ln b=3, to find each of the following
**
number in front of log is an exponent
when numbers multiply, logs add
when numbers divide, logs subtract
terms must be of the same base, e in this case
..
ln(ab)=lna+lnb=8+3=11
ln(b/a)=lnb-lna=3-8=-5
ln(b^4)=4 lnb=4*3=12
ln √(a)=1/2lna=1/2*8=4